find ordered pairs for 7y=4x+18 and y=-8x+54 please I cant seem to understand how to find solution by substitution????
Found 2 solutions by amoresroy, unlockmath:
Answer by amoresroy(361)

Equation 1:
7y=4x+18
Equation 2:
y=-8x+54
Substitute the value of y in equation 2 to equation 1
Equation 1:
7y=4x+18
7(-8x+54) = 4x + 18
7*(-8)x + 7*54 = 4x + 18
-56x + 378 = 4x + 18
378 - 18 = 4x + 56x
360 = 60x
x = 6
in equation 2
y = -8x + 54
y = -8(6) + 54
y = -48 +54
y = 6

Answer by unlockmath(1688)

Hello,
7y=4x+18 and y=-8x+54
Here's how we do it.
7(-8x+54)=4x+18
Rewritten as:
-56x+378=4x+18
Subtract 4x and 378 from both sides to get:
-60x=-360
Divide both sides by -60 to get:
x=6
Now we can find y by plugging 6 into either equation.
y=6
Make sense?
RJ
www.math-unlock.com